# Deep Neural Networks and Brain Alignment: Brain Encoding and Decoding (Survey)

### v2

@@ Abstract
Can we obtain insights about the brain using AI models? How is the information in deep learning models related to brain recordings? Can we improve AI models with the help of brain recordings? Such questions can be tackled by studying brain recordings like functional magnetic resonance imaging (fMRI). As a first step, the neuroscience community has contributed several large cognitive neuroscience datasets related to passive reading/listening/viewing of concept words, narratives, pictures, and movies. Encoding and decoding models using these datasets have also been proposed in the past two decades. These models serve as additional tools for basic cognitive science and neuroscience research. Encoding models aim at generating fMRI brain representations given a stimulus automatically. They have several practical applications in evaluating and diagnosing neurological conditions and thus may also help design therapies for brain damage. Decoding models solve the inverse problem of reconstructing the stimuli given the fMRI. They are useful for designing brain-machine or brain-computer interfaces. Inspired by the effectiveness of deep learning models for natural language processing, computer vision, and speech, several neural encoding and decoding models have been recently proposed. In this survey, we will first discuss popular representations of language, vision and speech stimuli, and present a summary of neuroscience datasets. Further, we will review popular deep learning based encoding and decoding architectures and note their benefits and limitations. Finally, we will conclude with a summary and discussion about future trends. Given the large amount of recently published work in the computational cognitive neuroscience (CCN) community, we believe that this survey enables an entry point for DNN researchers to diversify into CCN research.
Figure 1: Brain Encoding and Decoding: Datasets & Stimulus Representations. The plot of the encoding-decoding framework was derived from the study by Ivanova et al. (2022).
@@ 1 Introduction
The central aim of neuroscience is to unravel how the brain represents information and processes it to carry out various tasks (visual, linguistic, auditory, etc.). Two important models related to how brain represents information are, how external stimuli are represented in the form of neural responses (the encoding model) and how stimuli are recovered or reconstructed from the neuronal responses (the decoding model). The recent progress in deep neural networks in processing visual, auditory, linguistic, and multimodal stimuli makes one wonder if we could investigate these computational models and shed light on how the brain solves these problems. Thus, deep neural networks (DNN) may offer a computational medium to capture brain activities unprecedented complexity and richness of, leading to accurate encoding and decoding solutions. Previous surveys (Cao et al., 2021; Karamolegkou et al., 2023) have primarily focused on brain encoding and decoding studies for language stimuli. But recent attempts in cognitive neuroscience have focused on naturalistic and multimodal stimuli using DNNs. Hence, this survey systematically summarizes the latest encoding and decoding efforts on (i) how DNNs have begun to explain the underlying information processing in the brain for naturalistic stimuli of various modalities, (ii) the ways in which DNN models may be improved using the brain data, and (iii) the exploration of the shared underlying characteristics of both the systems.
@@
The survey aims to introduce the challenges in Computational Cognitive Neuroscience (CCN) to AI researchers familiar with recent advances in deep neural networks (DNNs). A good section of the DNN community is interested in neuroscience and psycholinguistics. Therefore, in this survey, we do not delve into architectural details and the learning procedures for DNNs rather, highlight how the advances in DNNs are used to address CCN problems. This enables an entry point for DNN researchers to diversify into CCN research.
Specifically, DNN researchers have the following advantages:
@@
    Using open-source brain datasets as evaluation benchmarks to improve DNN model capabilities further.
@@
    Training DNN models by incorporating brain recordings.
@@
    Developing better brain-computer interface (BCI) capabilities to decode brain patterns using advanced DNN models.
@@
    Clear exposition of various open source ecological stimuli datasets available and a curated GitHub repository for quick start of a study.
@@
    An accessible taxonomy of models and approaches.
@@
    A collection of open research problems in this fast-breaking research domain.
Brain encoding and decoding. Two main tools studied in cognitive neuroscience are brain encoding and brain decoding, as shown in Figure 1. Encoding is the process of learning the mapping $e$ from the stimuli $S$ to the neural activation $F$. The mapping can be learned using features engineering or deep neural networks. On the other hand, decoding constitutes learning mapping $d$, which predicts stimuli $S$ back from the brain activation $F$. However, in most cases, brain decoding aims to predict a stimulus representation $R$ rather than reconstructing $S$. In both cases, the first step is to learn a semantic representation $R$ of the stimuli $S$ at the train time. Next, a regression function $e:R\rightarrow F$ is trained for encoding. For decoding, a function $d:F\rightarrow R$ is trained. These functions $d$ and $e$ can then be used at test time to process new stimuli and brain activations, respectively. Ridge regression is the most popular choice for the functions $d$ and $e$.
To study the brain response to various modalities of stimuli, neuroscience researchers have curated several datasets. These datasets consist of stimuli and corresponding brain activity while participants were involved in interactions with the stimuli and optionally performing tasks such as language comprehension, visual and auditory processing, etc. Next, we discuss various techniques for obtaining the brain recordings and methods for representing stimuli.
Techniques for recording brain activations. Popular techniques for recording brain activations can be broadly classified into invasive and non-invasive techniques, as shown in Figure 2. Invasive techniques include single Micro-Electrode (ME), Micro-Electrode array (MEA), and Electro-Corticography (ECoG). The non-invasive recording techniques include functional magnetic resonance imaging (fMRI), Magneto-encephalography (MEG), Electro-encephalography (EEG) and Near-Infrared Spectroscopy (NIRS). Apart from the dimension of invasiveness, these techniques differ in their spatial resolution of neural recording and temporal resolution. fMRI recording enables data acquisition at high spatial but low temporal resolution. Hence, they are suitable for examining which parts of the brain handle critical functions. A typical whole brain fMRI acquisition takes 1-4 seconds to complete a scan. This is far slower than the speed at which humans can process language. On the other hand, both MEG and EEG have high temporal but low spatial resolution. They can preserve rich syntactic information (Hale et al., 2018) but cannot be used for source analysis. fNIRS offers a compromise option. The time resolution is better than fMRI, and spatial resolution is better than EEG. However, this spatial and temporal resolution balance may not compensate for the loss in both and its restriction in terms of only recording cortical activity but not from nuclei that are deeper in the brain, such as the basal ganglia, amygdala, hippocampus, etc. Further details on curation of brain recordings to specific brain regions are discussed in Section 4.
Stimulus representations. Neuroscience datasets contain stimuli across various modalities, including text, visual, audio, video, and other multimodal forms. Representations differ based on the modality. We briefly discuss extracting of stimulus representations from DNN models according to the following criteria: (1) Traditional and advanced models for text-based stimulus representations. (2) Image-based representations from deep vision models. (3) Extraction of low-level speech to Transformer-based speech-based auditory representations. (4) Finally, for multimodal stimulus representations, we explore early fusion and late fusion deep learning methods. Early fusion methods combine information across modalities at the initial processing stages, whereas late fusion combines it only at the end. Further details on different stimulus representation methods are discussed in Section 2.
@@
Figure 2: Overview of different brain–machine interfacing methods and their spatial and temporal resolution. Methods included: electroencephalography (EEG), magnetoencephalography (MEG), near-infrared spectroscopy (NIRS), functional magnetic resonance imaging (fMRI), electrocorticography (ECoG), microelectrode array (MEA) recordings and single microelectrode (ME) recordings. Adapted from (Van Gerven et al., 2009).
Figure 3: Representative Samples of Naturalistic Brain Datasets: (Left) Brain activity recorded when subjects are reading and listening to the same narrative (Deniz et al., 2019), and (Right) example naturalistic stimuli from various public repositories: BOLD5000 (Chang et al., 2019), ShortClips (Huth et al., 2022), Natural Scenes Dataset (NSD) (Allen et al., 2022) and Pereira dataset (Pereira et al., 2018).
Naturalistic neuroscience datasets. Several neuroscience datasets have been proposed across modalities (see Figure 3). These datasets differ in terms of the following criteria: (1) Method for recording activations: fMRI, EEG, MEG, etc. (2) Repetition time (TR), i.e. the sampling rate. (3) Characteristics of fixation points: location, color, shape. (4) Form of stimuli presentation: text, video, audio, images, or other multimodality. (5) Task that participant performs during recording sessions: question answering, property generation, rating quality, etc. (6) Time given to participants for the task, e.g., 1 minute to list given object properties. (7) Demography of participants: males or females, sighted or blind, etc. (8) Number of times the response to stimuli was recorded. (9) Natural language associated with the stimuli. We discuss details of proposed datasets in Section 3.
Evaluation of brain encoding and decoding methods. 2V2 accuracy and Pearson Correlation are two popularly used metrics for the evaluation of brain encoding models. On the other hand, brain decoding models are evaluated using metrics such as pairwise accuracy, rank accuracy, R2 score, and mean squared error. We discuss the detailed definitions of these metrics in Section 5.
Interpreting brain recordings through the robustness of DNN model representations. To interpret the stimulus representations obtained from DNN models and examine their impact on brain alignment, prior studies have proposed three different methods: variance partitioning (de Heer et al., 2017), the residual approach (Toneva et al., 2022a; Oota et al., 2024b), an indirect approach (Schrimpf et al., 2021; Goldstein et al., 2022), and the stacked regression approach (Lin et al., 2023). We discuss the details of each method in Section 6.3.
Computational Cognitive Neuroscience (CCN) research goals. CCN researchers have primarily focused on two main areas (Doerig et al., 2023).
@@
    Improving predictive accuracy. In this area, the work is around the following questions.
    • Compare feature sets: Which feature set provides the most faithful reflection of the neural representational space?
    • Test feature decodability: “Does neural data Y contain information about features X?”
    • Build accurate models of brain data: The aim is to enable the simulation of neuroscience experiments.
@@
    Interpretability. In this area, the work is around the following questions.
    • Examine individual features: Which contributes most to neural activity?
@@
    • Test correspondences between representational spaces: “CNNs vs ventral visual stream” or “Two text representations”.
@@
    • Interpret feature sets: Do features X, generated by a known process, accurately describe the space of neural responses Y? Do voxels respond to a single feature or exhibit mixed selectivity?
@@
We discuss some of these questions in Sections 6 and 7.
@@
Brain decoding models aim to understand what a subject is thinking, seeing, and perceiving by analyzing neural recordings. Over the past decades, the brain-computer interface (BCI) has made significant progress in decoding stimuli (language/images/speech) from the brain using non-invasive recordings. Like brain encoding literature, decoding literature focuses on studying a few important aspects: (1) In the context of language, how we compose the linguistic meaning from different stimuli such as text, images, videos, or speech by analyzing the evoked brain activity (Pereira et al., 2016; 2018). (2) Given brain activations corresponding to visual stimuli, how accurately can we decode a sentence representing the visual stimuli? (Nishimoto et al., 2011; Beliy et al., 2019) (3) How can we decode natural speech processing from non-invasive brain recordings using a single architecture and a data-driven approach? (Défossez et al., 2023) (4) How accurately can we reconstruct perceived natural images or decode their semantic contents from non-invasive recording data using popular deep learning models? (Takagi & Nishimoto, 2022). We discuss these details of decoding methods in Section 7.

### v3

@@ Abstract
Can artificial intelligence unlock the secrets of the human brain? How do the inner mechanisms of deep learning models relate to our neural circuits? Is it possible to enhance AI by tapping into the power of brain recordings? These captivating questions lie at the heart of an emerging field at the intersection of neuroscience and artificial intelligence. Our survey dives into this exciting domain, focusing on human brain recording studies and cutting-edge cognitive neuroscience datasets that capture brain activity during natural language processing, visual perception, and auditory experiences. We explore two fundamental approaches: encoding models, which attempt to generate brain activity patterns from sensory inputs; and decoding models, which aim to reconstruct our thoughts and perceptions from neural signals. These techniques not only promise breakthroughs in neurological diagnostics and brain-computer interfaces but also offer a window into the very nature of cognition. In this survey, we first discuss popular representations of language, vision, and speech stimuli, and present a summary of neuroscience datasets. We then review how the recent advances in deep learning transformed this field, by investigating the popular deep learning based encoding and decoding architectures, noting their benefits and limitations across different sensory modalities. From text to images, speech to videos, we investigate how these models capture the brain’s response to our complex, multimodal world. While our primary focus is on human studies, we also highlight the crucial role of animal models in advancing our understanding of neural mechanisms. Throughout, we mention the ethical implications of these powerful technologies, addressing concerns about privacy and cognitive liberty. We conclude with a summary and discussion of future trends in this rapidly evolving field. Given the large amount of recently published work in the computational cognitive neuroscience (CCN) community, we believe that this survey provides an invaluable entry point for deep neural network (DNN) researchers looking to diversify into CCN research, inviting them to join in unraveling the ultimate puzzle: the human brain.
@@ 1 Introduction
The central aim of neuroscience is to unravel how the brain represents information and processes it to carry out various tasks (visual, linguistic, auditory, etc.). Two critical paradigms in this field are encoding and decoding. The encoding model predicts neural responses from external stimuli, while the decoding model reconstructs or classifies stimuli from observed neural activity. Recent advancements in deep neural networks (DNNs) for processing visual, auditory, linguistic, and multimodal stimuli have raised intriguing questions about their potential to elucidate brain function. DNNs may offer a computational framework to capture the unprecedented complexity and richness of brain activities, potentially leading to more accurate encoding and decoding solutions. Building on this, previous surveys have addressed various aspects of brain encoding and decoding studies. For example, Cao et al. (2021) provided a general overview of mechanistic modeling in systems neuroscience, while Karamolegkou et al. (2023) focused on brain encoding and decoding studies for language stimuli. However, recent research in cognitive neuroscience has expanded to modeling for naturalistic and multimodal stimuli using DNNs. By integrating insights from studies involving diverse naturalistic stimuli, we can better understand the full potential of DNNs in modeling complex brain responses. Therefore, this survey aims to fill this gap by systematically summarizing the latest encoding and decoding efforts, focusing on:
  1. 1.
    How DNNs explain underlying information processing in the brain for naturalistic stimuli across various modalities.
  2. 2.
    Potential improvements to DNN models using brain data.
  3. 3.
    Exploration of shared characteristics between artificial and biological neural systems.
Figure 1: This figure summarizes overall encoding and decoding pipelines with different neuroimaging modalities (fMRI, MEG and EEG), Stimulus modalities (language, audio, visual, and multimodal), and Tasks (reading, listening, watching static images or videos, with or without audio). Further, the pipelines also incorporate stimulus representations obtained from different types of DNN, mapping the DNN and Brain representations via linear or non-linear models, and evaluation measures estimating the performance of encoding/decoding models. Visualization tools facilitate intuitive presentation of the results.
The general context of computational models of the brain function. The interdisciplinary field of Cognitive Computational Neuroscience (CCN) aims to combine findings from cognitive science, neuroscience, and computational modeling to understand how the brain represents and mediates cognitive processes. Computational models are pitched at three different scales: microscale, mesoscale, and macroscale. Microscale computational models are typically based on detailed biophysical modeling of the neurons and are aimed at deciphering the fundamental principles of neural computation (Dayan et al., 2003). Mesoscale models are designed to link the detailed dynamics of individual neurons and the large-scale organization of brain networks, typically using data from neuroimaging studies of the whole brain (Shine et al., 2021). However, macroscale computational models focus on understanding brain function at a larger systems-level, typically involving networks of brain regions and their interactions. The models discussed in the review belong to the mesoscale category. The cognitive architecture of the brain exhibits an intriguing interplay of structure and function that underpins human cognition, allowing the integration of sensory information (vision, auditory and other senses), perceptual processing, learning, memory, attention, motor processes, decision-making, emotions, and other higher-order cognitive processes. From the panoply of cognitive processes, the models reported here address only a small subset comprising perceptual processes related to vision and auditory domains, as well as behaviors related to reading, listening, speech, and language. Thus the new field of neuro-AI is still in its infancy, modeling the brain function in limited cognitive domains using the latest advances in deep learning for understanding how the brain encodes information and how to decode this back!
@@
This survey aims to introduce the challenges in Computational Cognitive Neuroscience (CCN) to AI researchers familiar with recent advances in DNNs. Rather than delving into architectural details and learning procedures for DNNs, we highlight how these advances are applied to CCN problems, providing an entry point for DNN researchers to diversify into CCN research.
Specifically, DNN researchers can benefit from this interdisciplinary approach in several ways:
@@
    Using open-source brain datasets as evaluation benchmarks to enhance DNN model capabilities.
@@
    Incorporating brain recordings in DNN training.
@@
    Developing advanced brain-computer interfaces (BCIs) for decoding brain patterns.
@@
    Accessing curated ecological stimuli datasets and a GitHub repository for efficient study initiation  [^1] https://github.com/subbareddy248/Awesome-Brain-Encoding--Decoding.
@@
    Understanding the taxonomy of CCN models and approaches.
@@
    Exploring open research problems in this rapidly evolving field.
Overall, Figure 1 provides an overview of how we move from human brain recordings to claims about computational principles using neural network models. Specifically, Figure 1 highlights various neuroimaging modalities, stimulus modalities, and tasks, incorporating stimulus representations derived from different types of DNNs. It also shows how DNN and brain representations are mapped through linear or non-linear models, the evaluation measures used to assess the performance of encoding/decoding models, and visualization tools that enable intuitive presentation of the results.
Brain encoding and decoding. Figure 1 illustrates the two main paradigms studied in cognitive neuroscience: brain encoding and decoding. Encoding involves learning a mapping $f_{e}$ from stimulus representations $\mathbf{X}$ to neural brain activation $\mathbf{Y}$, where $\mathbf{X}$ can be achieved through feature engineering or DNNs. Decoding, conversely, involves learning a mapping $f_{d}$ that predicts stimuli $\mathbf{X}$ from brain activation $\mathbf{Y}$. Often, brain decoding aims to predict a stimulus representation $\mathbf{X}$ rather than reconstructing stimuli $\mathcal{S}$ directly (see Table 1 for notation clarification). The process typically involves learning a semantic representation $\mathbf{X}$ of stimuli $\mathcal{S}$ during training, followed by training regression functions $f_{e}:\mathbf{X}\rightarrow\mathbf{Y}$ for encoding and $f_{d}:\mathbf{Y}\rightarrow\mathbf{X}$ for decoding. These functions can then be applied to new stimuli and brain activations during testing. Ridge regression is commonly used for both $f_{e}$ and $f_{d}$. To study brain responses to various stimulus modalities, neuroscientists have collected datasets comprising stimuli and corresponding brain activity. Participants interact with these stimuli, often performing tasks such as language comprehension or visual and auditory processing.
Table 1: Notations
Symbol Description
$\mathcal{S}$ Set of stimuli used in the experiment.
$\mathbf{X}$ Matrix representing stimulus data.
$\mathbf{Y}$ Matrix representing neural data.
$f_{e}$ encoding: function mapping from $\mathbf{X}\rightarrow\mathbf{Y}$
$f_{d}$ decoding: function mapping from $\mathbf{Y}\rightarrow\mathbf{X}$
$f_{e}(\mathbf{X})$ Matrix of predicted neural responses.
$f_{d}(\mathbf{Y})$ Matrix of predicted stimulus representation.
@@
Figure 2: Overview of different brain–machine interfacing methods and their spatial and temporal resolution. Methods included: electroencephalography (EEG), magnetoencephalography (MEG), near-infrared spectroscopy (NIRS), functional magnetic resonance imaging (fMRI), electrocorticography (ECoG), microelectrode array (MEA) recordings and single microelectrode (ME) recordings. Figure is adapted from (Van Gerven et al., 2009), and used with permission from the respective authors.
Techniques for recording brain activations. Popular techniques for recording brain activations can be broadly classified into invasive and non-invasive techniques, as illustrated in Figure 2. Invasive techniques include single Micro-Electrode (ME), Micro-Electrode array (MEA), and Electro-Corticography (ECoG). The non-invasive recording techniques include functional magnetic resonance imaging (fMRI), Magneto-encephalography (MEG), Electro-encephalography (EEG) and Near-Infrared Spectroscopy (NIRS). These techniques differ not only in their invasiveness but also in their spatial and temporal resolution. fMRI offers high spatial resolution but low temporal resolution, making it suitable for identifying which brain regions handle specific functions. A typical whole-brain fMRI scan takes 1-4 seconds, which is considerably slower than the speed of human cognitive processes. In contrast, MEG and EEG provide high temporal but low spatial resolution, capable of preserving rich syntactic information (Hale et al., 2018) but limited in source analysis. fNIRS offers a compromise, with better temporal resolution than fMRI and better spatial resolution than EEG. However, it is restricted to recording cortical activity and cannot capture signals from deeper brain structures such as the basal ganglia, amygdala, or hippocampus. Further details on the data collection of brain recordings for specific brain regions are discussed in Section 4.
Stimulus representations. Neuroscience datasets encompass stimuli across various modalities, including text, visual, audio, video, and other multimodal forms. We briefly discuss the extraction of stimulus representations from DNN models according to the following criteria:
@@
    Traditional and advanced models for text-based stimulus representations.
@@
    Image-based representations from deep vision models.
  3. 3.
    Extraction of low-level speech to Transformer-based speech auditory representations.
  4. 4.
    Multimodal stimulus representations, exploring early fusion (combining information across modalities at initial processing stages) and late fusion (combining at the end) deep learning methods.
Further details on different stimulus representation methods are discussed in Section 2.
Figure 3: Representative Samples of Naturalistic Brain Datasets. (Left) Comparison of brain activity patterns recorded during reading and listening to the same narrative, illustrating modality-specific and shared neural responses (Deniz et al., 2019). (Right) Examples of diverse naturalistic stimuli used in various public neuroimaging repositories: complex visual scenes from BOLD5000 (Chang et al., 2019), video frames from ShortClips (Huth et al., 2022), natural images from the Natural Scenes Dataset (NSD) (Allen et al., 2022), and multimodal stimuli (text and images) from the Pereira dataset (Pereira et al., 2018). Adapted with permission from the respective authors.
Naturalistic neuroscience datasets involving human participants. Numerous neuroscience datasets have been proposed across modalities (see Figure 3). These datasets vary in terms of: (1) Method for recording activations (fMRI, EEG, MEG, etc.); (2) Repetition time (TR), i.e., the sampling rate; (3) Characteristics of fixation points (location, color, shape); (4) Form of stimuli presentation (text, video, audio, images, or multimodal); (5) Task (question answering, property generation, rating quality, etc.) performed by participants during recording sessions; (6) Time allocated for task completion, e.g., 1 minute to list given object properties; (7) Participant demographics, e.g. males or females, sighted or blind; (8) Number of repetitions for stimuli response recording; (9) Natural language associated with the stimuli. We discuss details of these datasets in Section 3.
From Non-Human Models to Human Brain-Machine Alignment. While this review focuses on brain-machine alignment using human neuroimaging datasets, it is important to acknowledge the substantial contributions made by research involving non-human subjects. Neuro-AI and cognitive-AI methods have made significant strides in modeling the primate visual system and hippocampal navigation system by leveraging advanced machine learning and deep learning techniques to mimic neural processes. Since high-resolution data related to hierarchically organized maps of stimulus properties such as orientation, spatial frequency, and color, are not available for human visual system, measurements from the primate (macaque V1) data have been used to design and validate deep learning models of visual function (Margalit et al., 2024). Several non-human primate datasets, like the primate microelectrode recording dataset proposed in Majaj et al. (2015), are highly valuable and widely used in vision research. The other area where tremendous progress has been made is in the hippocampal memory formation and spatial navigation using rodent electrophysiology data. These efforts led to newer understanding of how place cells and grid cells in the hippocampus and entorhinal cortex encode spatial information and brain’s ability to build cognitive maps for spatial navigation (Bellmund et al., 2018). The emerging insights of brain encoding and decoding enabled successful design and testing of brain-machine interfaces for primates (Lebedev & Nicolelis, 2017). Successful application of deep neural networks in mimicking primate perceptual functions have led to formulation of newer frameworks for systems neuroscience research (Richards et al., 2019). In contrast, this review focuses exclusively on neuroscience datasets involving human participants, particularly those with non-invasive recordings such as fMRI, MEG, and EEG. Given the scope of our current review paper, our focus is limited to human brain datasets with an emphasis on the language function that is difficult, if impossible, to investigate in rodents and non-human primates.
Evaluation of brain encoding and decoding methods. For brain encoding models, 2-versus-2 (2V2) accuracy, Pearson Correlation and R2 score are commonly used evaluation metrics. Brain decoding models are typically evaluated using metrics such as pairwise accuracy, rank accuracy, R2 score, and mean squared error. Additionally, representational similarity metrics frequently used in brain encoding and decoding studies include representational dissimilarity matrices (RDM), centered kernel alignment (CKA), and normalized distribution similarity (NDS). Detailed definitions of these metrics are provided in Section 5.
Interpreting brain recordings through DNN model representations. To interpret stimulus representations obtained from DNN models and examine their impact on brain alignment, prior studies have proposed several methods: variance partitioning (de Heer et al., 2017), the residual approach (Toneva et al., 2022a; Oota et al., 2024b), an indirect approach (Schrimpf et al., 2021; Goldstein et al., 2022), and the stacked regression approach (Lin et al., 2023). We discuss these methods in detail in Section 6.3.
Computational Cognitive Neuroscience (CCN) research goals. CCN researchers have primarily focused on two main areas (Doerig et al., 2023):
  1. 1.
    Improving predictive accuracy, addressing questions such as:
    • Which feature set best reflects the neural representational space?
    • Does neural data Y contain information about features X?
    • How can we build accurate models of brain data to enable neuroscience experiment simulation?
  2. 2.
    Enhancing interpretability, exploring questions like:
    • Which features contribute most to neural activity?
@@
    • How do representational spaces correspond, for example, between CNNs and the ventral visual stream, or among text representations?
@@
    • Do features X, generated by a known process, accurately describe the space of neural responses Y?
    • Do voxels respond to single features or exhibit mixed selectivity?
@@
Overall, the CCN community tries to understand the brain in computational terms which also involves falsifying existing models, and making new (model-guided) predictions (Tuckute et al., 2024). We discuss some of these questions in Sections 6 and 7.
@@
Brain decoding models aim to understand what a subject is thinking, seeing, and perceiving by analyzing neural recordings. Over the past decades, the brain-computer interface (BCI) has made significant progress in decoding stimuli (language/images/speech) from the brain using non-invasive recordings. Like brain encoding literature, decoding literature focuses on studying a few important aspects: (1) In the context of language, how do we compose the linguistic meaning from different stimuli such as text, images, videos, or speech by analyzing the evoked brain activity (Pereira et al., 2016; 2018). (2) Given brain activations corresponding to visual stimuli, how accurately can we decode a sentence representing the visual stimuli? (Nishimoto et al., 2011; Beliy et al., 2019) (3) How can we decode natural speech processing from non-invasive brain recordings using a single architecture and a data-driven approach? (Défossez et al., 2023) (4) How accurately can we reconstruct perceived natural images or decode their semantic contents from non-invasive recording data using popular deep learning models? (Takagi & Nishimoto, 2022). We discuss these details of decoding methods in Section 7.